Problem 14-4 Bond amortization schedule [LO14-2) On January 1, 2018, Tennessee Harvester Corporation issued debenture bonds that pay interest semlannually on June 30 and December 31. Portions of the bond amortization schedule appear below 280,000293,611 280,000 280,000295,293 280,000 296,211 280,000297,18 280,000 298,215 13,611 14,428 15,293 16,211 17,184 18,215 4,907,129 294,428 4,953,061 4,970,245 4.988,46o 38 39 40 280,000397,547 280,000404, 600 280,000 412,072 117,547 124,600 132,072 6,867,928 7,000,000 1. What is the face amount of the bonds? 2. What is the initial selling price of the bonds? 3. What is the term to maturity in years? 4. Interest is determined by what approach? 5. What is the stated annual interest rate? 6. What is the effective annual interest rate? 7. What is the total cash interest paid over the term to maturity? 8. What is the total effective interest expense recorded over the term to maturity? < Prey 7 of 10EE Next >
7. What is the total cash interest paid over the term to maturity? 8. What is the total effective interest expense recorded over the term to ma 1. Face amount 2. Initial selling price 3. Term to maturity 4. Interest is determined by what approach? 5. Annual interest rate 6. Effective annual interest rate 7. Total cash interest paid years Effective interest expense
Answer #1
A B C D E F G H I J
2
3 1)
4
5 Face amount will be the outstanding Balance at the maturity i.e. at payment 40.
6
7 Thus face value is $7,000,000
8
9 2)
10 Initial selling price of the bond will be the outstanding balance with zero payment
11 i.e. 4,893,518.00
12
13 Hence initial selling price is $4,890,518.00
14
15 3)
16
17 Since the bond is paying the coupon semi-annually, therefore
18 term of maturity in years will be the half of the total number of payments.
19
20 Since total number of payments is 40 therefore term of maturity in years will be 20 years.
21
22 Hence term of maturity in years will be 20 years
23
24 4)
25
26 Effective interest rate method has been used to calculated the interest.
27
28 5)
29
30 Face Value $7,000,000
31 Coupon Amount $280,000
32
33 Semi-annual interest 4.00% =D31/D30
34
35 Hence stated annual interest rate 8.0%
36
37 6)
38 For the first period
39 Interest rate =Effective interest / Oustanding Balance before 1st payment
40 =293,611 / 4,893,518
41 6.00%
42
43 Hence the interest rate is 6.00%
44
45 7)
46
47 Total interest paid =Payment Amount*Number of payments
48 =280,000*40
49 $11,200,000
50
51 Hence total interest paid is $11,200,000
52
53 8)
54
55 Total effective interest expense recorded =Total interest Paid + Total Discount Amount amortized
56 =11,200,000+(7,000,000-4,893,518)
57 $13,306,482
58
59 Hence total effective interest expense is $13,306,482
60